Lebron is 5 years older than Kuz. The product of their ages is 21 years less than 15 times the sum of their ages. How old is Kuz?
Answer by ewatrrr(24785)

 
Hi
Let Kuz' age be x, the Lebron's age is (x+5)
The product of their ages is 21 years less than 15 times the sum of their ages
Write as You Read:
x(x+5) = 15(x + (x+5)) -21
Simply and solve:
x^2 + 5x = 15(2x+5)-21
x^2 + 5x = 30x + 75 - 21
x^2 -25x - 54 = 0
x = 27 (tossing out negative value)
Kuz is 27 years old
Wish You the Best in your Studies.